# SLIM: Let LLM Learn More and Forget Less with Soft LoRA and Identity Mixture

### v1

@@ 3.5 Summarising the proposed SLIM
To learn the downstream tasks while mitigating catastrophic forgetting, the proposed SLIM modifies the LoRA architecture on the following sides:
@@
    We introduce weight yielding with sliding clustering to correct the routing weight and encourage the model to predict downstream tasks with the trained adapters, and the general tasks without them.
@@
For a given input, we first assign it to its nearest cluster center and calculate the distance in between. During training, the centers and the distributions of the clusters are updated according to the assigned samples. Then the router predicts the routing weight of the input. The routing weight is then yielded according to the distance, as depicted in Eq. 9, and then normalized to the final routing weight as in Eq. 10. We activate the mixed experts including identity layers and LoRA adapters according to the routing weight, and utilizing the dynamic merging to obtain the final output. The whole process could be summarized as
$\begin{array}[]{l}y=Wx+b+\sum_{i}\hat{w}_{i}f_{i}(x),\\ f_{i}(x)=\left\{\begin{array}[]{c}x,i\leq K\\ B^{\prime}_{i}A^{\prime}_{i}x,otherwise\end{array},\right.\end{array}$ (18)
@@ 4.1 Implementation details
We evaluate the proposed SLIM on the Openchat-8B model Wang et al. (2024), an extension of Llama3-8B-Instruct. We conduct two different training settings: single dataset setting (SDS) and multi-dataset setting (MDS). For SDS, we train the model with a single downstream task. For MDS, we mix the datasets for model training. MDS include OBQA Mihaylov et al. (2018), SIQA Sap et al. (2019) and BOOLQ Clark et al. (2019) datasets. For SDS, we involve CSQA Talmor et al. (2019), Hellaswag Zellers et al. (2019), Winogrande Sakaguchi et al. (2021), ARC-e and ARC-c Clark et al. (2018) datasets. The fine-tuned models under MDS are evaluated on the general tasks including MMLU Hendrycks et al. (2021), GSM8K Cobbe et al. (2021), and PIQA Bisk et al. (2020) to measure the catastrophic forgetting of the PEFT approaches. We fine-tune the model 2 epochs for all tasks, on a single NVIDIA-A100 80G GPU, the batch size is set to 16.
@@ 4.2 Comparison with SOTA approaches
First, we compare the proposed approach with other widely used PEFT approaches. LoRAMoE, MixLoRA, MoLA are set to $rank=16,N_{\text{experts}}=8$, and SLIM holds $N_{\text{experts}}=10$ with 2 identity layers as experts, contains the same number of parameters. For a fair comparison, we set the $rank$ of LoRA and DoRA to 128. The results are shown in Tab. 1. The result demonstrates that the proposed approach achieves comparable performance on the downstream tasks, and achieves the best performance on multiple datasets. The models fine-tuned under the MDS setting are further evaluated on the general datasets, and the results are shown in Tab. 2. Compared with the other PEFT approaches, the proposed SLIM method significantly alleviates the loss of generalization capacity of the fine-tuned model, without any extra data replay.
@@ 4.3 Ablation study
We first validate the effectiveness of the proposed modules, and the results are shown in Tab. 3. The baseline is the MixLoRA approach. Simply removing the LoRA adapters for the attention layers has a minor effect on the MDS dataset, but reduces the forgetting by 2.29%. The insertion of identity layers reduces the forgetting by 0.96%, but resulting in a performance drop on the MDS. The weight yielding further boosts the model by 2.59% compared with the baseline model without LoRA for attention layers. The dynamic merging significantly boots the model on MMLU by 5.54%, demonstrating its effectiveness for forgetting alleviation. Surprisingly, dynamic merging also improves the model performance on the MDS datasets, which might be because the general capability could benefit some downstream tasks.
@@
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ 79.10 60.64
@@
To validate the efficiency and effectiveness of the proposed fast merging approach, we conduct an experiment on the MDS and the MMLU dataset, and the results are demonstrated in Tab. 4. The fast implementation of the dynamic merging has a slight influence on the model performance, while dramatically reducing the time cost.
@@
As the masking ratio controls the contribution of the LoRA experts, we conduct experiments to measure this influence. As depicted in Eq. 14, we sample rows from the low-rank matrices.It is worth noting that the masking rate $p$ is quadratic to the masking rate of the rows and columns $p_{rc}$. Specifically, $p_{rc}=1-\sqrt{p}$. For convenience, we validate the masking rate of rows and columns, instead of directly validating the masking rate.
The experiment results are shown in Fig. 4. If the masking ratio is too large, the LoRA experts will be eliminated from the model, resulting in a significant performance drop in the downstream tasks. On the contrary, a small masking ratio can not inhibit catastrophic forgetting, therefore resulting in a degenerated performance for MMLU. We find that utilizing a balanced masking rate $p$ (e.g., 0.5) boosts both downstream tasks while maintaining the generalization of the model, which may be because the downstream task could somewhat benefit from the general capability of the instruction tuning model.
@@
Figure 6: The validation of the proposed SLIM on MDS (downstream) and MMLU (general) with different number of LoRA experts.
@@ 5 Conclusion
In this work, we propose SLIM, an MoE parameter-efficient fine-tuning approach. In addition to implementing the LoRA experts, SLIM also adopts identity layers as experts and dynamically routes between the LoRA adapters and identity layers. We also propose to yield the routing weights according to the distances in the latent space. We further propose dynamic merging that formulates PEFT to model merging, which effectively combines the downstream capabilities with the original capabilities of the pre-trained model. The proposed approach achieves comparable performance to the state-of-the-art PEFT methods, with much fewer capability loss.
@@ 6.2 Difficulty for implementation
Many implementation and acceleration frameworks have been proposed for LLMs, such as vLLM vLLM Team and ollama ollama Team . However, despite the LoRA adapters could be re-parameterized to the base model, and vLLM support implementing multiple LoRA adapters to the same base model, it is still a challenge to implement LLMs with mixed heterogeneous adapters (LoRA adapters and identity layers). We are working on the acceleration engine for the MoE of LoRA adapters for its further utilization.

### v2

@@ 3.5 Summarising the proposed SLIM
To adapt to the downstream tasks while mitigating catastrophic forgetting, the proposed SLIM modifies the LoRA architecture on the following sides:
@@
    We introduce weight yielding with sliding clustering to correct the routing weight and encourage the model to predict downstream tasks with the fine-tuned adapters, and the general tasks without them.
@@
For a given input, we first assign it to its nearest cluster center and calculate the distance in between. During training, the centers and the distributions of the clusters are updated according to the assigned samples. Then the router predicts the routing weight of the input. The routing weight is then yielded according to the distance, as depicted in Eq. 10, and then normalized to the final routing weight as in Eq. 11. We activate the mixed experts including identity layers and LoRA adapters according to the yielded routing weight, and utilizing the dynamic merging to obtain the final output. The overall process could be summarized as:
$\begin{array}[]{l}y=Wx+b+\frac{1}{Z}\sum_{i}\hat{w}_{i}f_{i}(x),\\ f_{i}(x)=\left\{\begin{array}[]{c}\textbf{0},i\leq K\\ B^{\prime}_{i}A^{\prime}_{i}x,otherwise\end{array},\right.\end{array}$ (19)
@@ 4.1 Implementation details
We evaluate the proposed SLIM on the OpenChat-8B model Wang et al. (2024), an extension of Llama3-8B-Instruct. We conduct two different training settings: single dataset setting (SDS) and multi-dataset setting (MDS). For SDS, we train the model with a single downstream task. For MDS, we mix the datasets for the model fine-tuning. MDS include OBQA Mihaylov et al. (2018), SIQA Sap et al. (2019) and BOOLQ Clark et al. (2019) datasets. For SDS, we involve CSQA Talmor et al. (2019), Hellaswag Zellers et al. (2019), Winogrande Sakaguchi et al. (2021), ARC-e and ARC-c Clark et al. (2018) datasets. The fine-tuned models under MDS are evaluated on the general tasks including MMLU Hendrycks et al. (2021), GSM8K Cobbe et al. (2021), and PIQA Bisk et al. (2020) to measure the catastrophic forgetting of the PEFT approaches. We fine-tune the model 2 epochs for all tasks, on a single NVIDIA-A100 80G GPU, the batch size is set to 16.
@@ 4.2 Comparison with SOTA approaches
First, we compare the proposed approach with other widely used PEFT approaches. Following Li et al. (2024), LoRAMoE Dou et al. (2024), MixLoRA Li et al. (2024), MoLA Gao et al. (2024) are set to $rank=16$, $N_{\text{experts}}=8$, and SLIM holds $N_{\text{experts}}=10$ with $K=2$ identity layers as experts, contains the same number of parameters. For a fair comparison, we set the $rank$ of LoRA and DoRA to 128. The results are shown in Tab. 1. The proposed approach achieves comparable performance on the downstream tasks and the best performance on multiple datasets. The models fine-tuned under the MDS setting are further evaluated on the general tasks, and the results are shown in Tab. 2. Compared with the other PEFT approaches, the proposed SLIM method significantly alleviates the loss of generalization capacity of the fine-tuned model, without any extra data replay.
@@ 4.3 Ablation study
We first validate the effectiveness of the proposed modules, and the results are shown in Tab. 3. The baseline is the MixLoRA approach. Simply removing the LoRA adapters for the attention layers has a minor effect on the MDS dataset, but reduces the forgetting by 2.29%. The insertion of identity layers reduces the forgetting by 0.96% but results in a performance drop on the MDS. Combining weight yielding with identity layers boosts the model by 1.73% on MDS, and alleviates forgetting by 2.59%. This indicates some MDS tasks could benefit from the general capabilities, and the weight yielding further enhances the mixed utilization of world and specific knowledge. The dynamic merging significantly boots the model on MMLU by 5.54%, demonstrating its effectiveness for forgetting alleviation. Although mixing identity layers and LoRA adapters benefit downstream tasks, it does not imply that fine-tuning is ineffective, as demonstrated by the following experiments.
@@
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ 79.53 60.29
@@
To validate the efficiency and effectiveness of the proposed fast implementation of dynamic merging, we conduct an experiment on the MDS and the MMLU dataset, and the results are demonstrated in Tab. 4. The fast implementation of the dynamic merging has a slight influence on the model performance, while dramatically reducing the time cost.
@@
As the masking ratio controls the contribution of the LoRA experts, we conduct experiments to measure this influence. As depicted in Eq. 15, we sample rows from the low-rank matrices. It is worth noting that the masking rate $p$ is quadratic to the masking rate of the rows and columns $p_{rc}$. Specifically, $p=1-(1-p_{rc})^{2}$. For convenience, we validate the masking rate of rows and columns, instead of directly validating the masking rate. The experiment results are shown in Fig. 4. If the masking ratio is too large, the LoRA experts will be eliminated from the model, resulting in a significant performance drop in the downstream tasks. On the contrary, a small masking ratio can not inhibit catastrophic forgetting, therefore resulting in a degenerated performance for MMLU. We find that utilizing a balanced masking rate $p$ (0.4$\sim$0.6) boosts both downstream tasks while maintaining the generalization of the model, which indicates the downstream task could benefit from the general capability of the instruction tuning model.
@@
Figure 6: The validation of the proposed SLIM on MDS (downstream) and MMLU (general) with different numbers of LoRA experts.
@@ 5 Conclusion
In this work, we propose SLIM, an MoE parameter-efficient fine-tuning approach. In addition to implementing the LoRA adapters, SLIM adopts identity layers as experts which could be activated as information "highways" to bypass the LoRA adapters for general capacity preservation. We propose weight yielding with sliding clustering to modify the routing weight according to the consistency between the input sample and the downstream distribution to take advantage of these “highways”. We further introduce dynamic merging that formulates PEFT to model merging, which effectively combines the downstream capabilities with the general capabilities of the base model. SLIM achieves comparable performance to state-of-the-art PEFT methods, with much less general capability loss.
@@ 6.2 Difficulty for implementation
Many implementation and acceleration frameworks have been proposed for LLM, such as vLLM vLLM Team and ollama ollama Team . However, despite the LoRA adapters could be re-parameterized to the base model, and vLLM support implementing multiple LoRA adapters to the same base model, it is still a challenge to implement LLM with mixed heterogeneous experts (LoRA adapters and identity layers). We are working on the acceleration engine for the MoE of LoRA adapters for its further utilization.
Appendix A Appendix
A.1 Case study on GSM8K for identity layers
To further demonstrate the effectiveness of the identity layers that skip the LoRA adapters, we conduct a case study on the GSM8K dataset. For the baseline, we utilize the MixLoRA model without LoRA adapters for the attention layers, denoted as “w/o iden”, and the proposed method is denoted as “w/ iden”. Note that the dynamic merging is discarded for a fair comparison. We utilize the red to mark the first mistake. The results are listed as follows:
Question 1: Toula went to the bakery and bought various types of pastries. She bought 3 dozen donuts which cost \$68 per dozen, 2 dozen mini cupcakes which cost \$80 per dozen, and 6 dozen mini cheesecakes for \$55 per dozen. How much was the total cost?
w/ iden: Toula bought 3 dozen donuts which cost \$68 per dozen, so she spent 3 * \$68 = \$204 on donuts. She bought 2 dozen mini cupcakes which cost \$80 per dozen, so she spent 2 * \$80 = \$160 on cupcakes. She bought 6 dozen mini cheesecakes for \$55 per dozen, so she spent 6 * $55=$330 on cheesecakes. The total cost was \$204 + \$160 + \$330 = \$694. The answer is 694
w/o iden: The cost of the donuts was 3 x \$68 = \$204. The cost of the mini cupcakes was 2 x \$80 = \$160. The cost of the mini cheesecakes was 6 x \$55 = \$330. The total cost was \$204 + \$160 + \$330 = \$694.
Question 2: Marie ordered one chicken meal that costs \$12, 5 packs of milk that costs \$3 each, 4 apples that cost \$1.50 each, and some boxes of pizza. Marie paid a total of \$50. How many boxes of pizza did Marie order if each box costs \$8.50?
w/ iden: Marie paid \$12 for the chicken meal, \$3 x 5 = \$15 for the milk, \$1.50 x 4 = \$6 for the apples, and \$50 - \$12 - \$15 - \$6 = \$17 for the pizza. \$17 / \$8.50 = 2 boxes of pizza The answer is 2
w/o iden: Marie paid \$12 for a chicken meal, \$3 for 5 packs of milk, \$1.50 for 4 apples, and \$50 in total. So, \$50 - \$12 - \$3 - \$1.50 = \$33.50 was spent on the boxes of pizza. Therefore, \$33.50 / \$8.50 per box = 4 boxes of pizza were ordered. The answer is 4
Question 3: Gloria is shoe shopping when she comes across a pair of boots that fit her shoe budget. However, she has to choose between the boots and two pairs of high heels that together cost five dollars less than the boots. If one pair of heels costs \$33 and the other costs twice as much, how many dollars are the boots?
w/ iden: The cost of the two pairs of heels is \$33 + \$33 * 2 = \$99. The cost of the boots is \$99 + \$5 = \$104. The answer is 104
w/o iden: The first pair of heels cost \$33 and the second pair cost twice as much, so the second pair cost 33 * 2 = \$66. The two pairs of heels together cost 33 + 66 = \$99. The boots cost five dollars less than the two pairs of heels, so the boots cost 99 - 5 = \$94. The answer is 94
Question 4: A treasure hunter found a buried treasure chest filled with gems. There were 175 diamonds, 35 fewer rubies than diamonds, and twice the number of emeralds than the rubies. How many of the gems were there in the chest?
w/ iden: The number of rubies is 175 diamonds - 35 rubies = 140 rubies. The number of emeralds is 140 rubies x 2 = 280 emeralds. The total number of gems is 175 diamonds + 140 rubies + 280 emeralds = 595 gems. The answer is 595
w/o iden: The number of rubies was 35 fewer than the diamonds, so there were 175-35 = 140 rubies. The number of emeralds was twice the number of rubies, so there were 140*2 = 280 emeralds. The total number of gems in the chest was 175 diamonds + 140 rubies + 280 emeralds = 495 gems. The answer is 495
Question 5: Vincent can buy flowers in packages of 3 for \$2.50 or in packages of 2 for \$1. How much money does he save by buying 18 flowers at the better price?
w/ iden: Vincent can buy 3 flowers for \$2.50, so the cost per flower is \$2.50 / 3 = \$0.8333. He can buy 2 flowers for \$1, so the cost per flower is \$1 / 2 = \$0.50. The difference in cost per flower is \$0.8333 - \$0.50 = \$0.3333. Vincent wants to buy 18 flowers, so he saves 18 * \$0.3333 = \$6.00 by buying the flowers at the better price. The answer is 6
w/o iden: Vincent can buy 6 flowers for \$2.50 and 9 flowers for \$1. Therefore, he can buy 15 flowers for \$3.50. If he buys 18 flowers at the better price, he will spend \$3.50 + \$1 = \$4.50. Thus, he saves \$4.50 - \$3.50 = \$1.00. The answer is 1
For the first question, the baseline model failed to follow the formulations of the examples (summarize the output with “The answer is”). For the rest of the questions, the baseline misunderstands the question or calculates incorrectly. This result indicates that the skipping of LoRA adapters via identity layers enhances the instruction-follow ability of the fine-tuned model, and mitigates the degeneration of instruction understanding and calculation.
Fig. A1.  The standard dynamic merging pipeline. Compared with the fast implementation, it needs to obtain the full-size weight matrix $BA$.
A.2 Standard dynamic merging pipeline
In Fig. A1, we demonstrate the standard dynamic merging pipeline. The standard dynamic merging needs to first calculate and mask $BA$, then multiply the result with the hidden state. Since $r<<\text{min}(C_{\text{in}},C_{\text{out}})$, calculate $M\otimes(BA)x$ is not efficient. On the contrary, the fast implementation calculates $B^{\prime}(A^{\prime}x)$, therefore reducing the computation cost.
A.3 Details of the involved datasets
OBQA
OpenBookQA is a question-answering dataset modeled after open book exams for assessing human understanding of a subject. It consists of 5,957 multiple-choice elementary-level science questions (4,957 train, 500 dev, 500 test).
SIQA
Social Interaction QA (SIQA) is a question-answering benchmark for testing social commonsense intelligence. Social IQa focuses on reasoning about people’s actions and their social implications. It consists of 37,588 QA pairs ( 33,410 train, 1,954 dev, 2,224 test).
BOOLQ
BoolQ is a binary QA dataset containing 15942 examples, of which 9.4k for the train set, 3.2k for the dev set, and 3.2k for the test set.
HellaSwag
HellaSwag is a sentence-completion multi-choice dataset, that contains 50k samples, in which 40k for training, 10k for test.
WinoGrande
WinoGrande is a common-sense reasoning dataset, including 43,972 samples (40,938 for training, 1,267 for development, and 1,767 for testing).
CSQA
The Commonsense QA (CSQA) is a dataset for commonsense question-answering tasks. The dataset consists of 12,247 questions with 5 choices each and is split in a (80/10/10) distribution.
ARC-c and ARC-e
ARC is a commonsense reasoning dataset, consisting of ARC-c and ARC-e. ARC-c contains 2590 samples, 1172 for the test, and ARC-e contains 5197 samples, 2376 for the test.
GSM8K
The GSM8K dataset is a mathematics dataset that contains 8.5K samples, of which 1.3k are for testing.
MMLU
MMLU is a benchmark dataset that contains 57 subjects from STEM, containing 14079 test samples.
PIQA
The PIQA (Physical Interaction Question Answering) dataset is designed to evaluate a model’s understanding of physical commonsense reasoning. It contains over 16,000 training samples and about 2k and 3k samples are held out as the dev set and test set.
All the mentioned datasets are open-sourced and allow academic use. We report results for the test set when the ground truth is available, otherwise, use the dev set.